# HTMNet: A Hybrid Network with Transformer-Mamba Bottleneck Multimodal Fusion for Transparent and Reflective Objects Depth Completion

### v1

@@ Abstract
Transparent and reflective objects pose significant challenges for depth sensors, resulting in incomplete depth information that adversely affects downstream robotic perception and manipulation tasks. To address this issue, we propose HTMNet, a novel hybrid model integrating Transformer, CNN, and Mamba architectures. The encoder is constructed based on a dual-branch Transformer-CNN framework, while the multi-scale fusion module leverages a Transformer-Mamba architecture, which also serves as the foundation for the decoder design. We introduce a novel multimodal fusion module grounded in self-attention mechanisms and state space models, marking the first application of the Mamba architecture in the field of transparent object depth completion and revealing its promising potential. Additionally, we design an innovative multi-scale fusion module for the decoder that combines channel attention, spatial attention, and multi-scale feature extraction techniques to effectively integrate multi-scale features through a down-fusion strategy. Extensive evaluations on multiple public datasets demonstrate that our model achieves state-of-the-art(SOTA) performance, validating the effectiveness of our approach.
@@ I INTRODUCTION
Models based on Transformer and CNN have been widely applied in the field of transparent object depth completion. To the best of our knowledge, however, there are currently no reported applications of the Mamba architecture in this domain. To address this long-standing gap, we conduct our research to explore the potential of the Mamba architecture in this context, aiming to open new research avenues for the community.
@@
  • We conduct extensive evaluations on multiple public datasets for transparent object depth completion, achieving state-of-the-art(SOTA) performance.
@@ II Related Works
II-A Depth Completion for Single View Transparent Objects
@@
Depth completion, as a fundamental task in computer vision, typically uses raw depth and RGB images to recover missing depth values in depth maps[13][14][15][16]. Due to the convenience of single-view acquisition and rapid data capture, depth completion for transparent and reflective objects under monocular settings has become a research hotspot. Before end-to-end models were applied to depth completion, two approaches were adopted for transparent object depth completion. The first relies on external assumptions, such as specialized sensing devices[17] and predefined backgrounds[18]. However, these methods are unsuitable for random environments due to their dependency on external assumptions. To overcome these limitations, a second approach using global optimization algorithms emerged[19]. Studies in[20],[21]utilized sensor-captured RGB-D images as input for depth completion, eliminating the need for external assumptions. However, they rely on surface normal and edge predictions, a limitation later addressed by end-to-end models. In[22], a generative adversarial network (GAN) was employed to generate depth maps for transparent object reconstruction. In[23], a multi-scale depth network was proposed, combining RGB and depth images as input. Although both ToDE-Trans[24] and swinDRNet[5] adopt Swin Transformers, they differ significantly in architecture. Specifically, swinDRNet is a dual-stream fusion network that integrates raw depth information with predicted depth through confidence maps, while ToDE-Trans follows a classical encoder-decoder structure. In[25], a decoder with multiple cascaded ”refinement modules” was designed. In[26], to enable cross-level feature fusion and better leverage low-level depth features, researchers introduced a skip-connected fusion branch within a U-Net framework.
II-B CNN-Transformer Hybrid Architecture for Vision Tasks
In visual recognition tasks, Vision Transformer (ViT) models have achieved remarkable success due to their superior performance[27]. However, the multi-head self-attention (MHA) mechanism not only incurs high computational costs but also struggles to capture inductive biases for local relationships. To address this issue, numerous studies have explored novel approaches to integrate convolutional mechanisms with self-attention. For instance, Li et al.[28]introduced MSA at the front end of the model while employing CNNs at the backend, a design that contrasts sharply with the method in[29], where CNNs first extract local features and MSA then learns long-range dependencies. In certain specialized tasks, the parallel use of CNNs and Transformers has gradually become a common strategy. Taking[30]as an example, researchers typically extract features separately through CNN and Transformer branches, followed by feature fusion at later stages. Additionally, in[31]and[32], convolutional and attention mechanisms are jointly utilized to construct modules, which are iteratively applied throughout the network. Despite the diversity of these fusion strategies, their core objective remains consistent: enabling models to simultaneously learn global and local information.

### v2

@@ Abstract
Transparent and reflective objects pose significant challenges for depth sensors, resulting in incomplete depth information that adversely affects downstream robotic perception and manipulation tasks. To address this issue, we propose HTMNet, a novel hybrid model integrating Transformer, CNN, and Mamba architectures. The encoder is based on a dual-branch CNN-Transformer framework, the bottleneck fusion module adopts a Transformer-Mamba architecture, and the decoder is built upon a multi-scale fusion module. We introduce a novel multimodal fusion module grounded in self-attention mechanisms and state space models, marking the first application of the Mamba architecture in the field of transparent object depth completion and revealing its promising potential. Additionally, we design an innovative multi-scale fusion module for the decoder that combines channel attention, spatial attention, and multi-scale feature extraction techniques to effectively integrate multi-scale features through a down-fusion strategy. Extensive evaluations on multiple public datasets demonstrate that our model achieves state-of-the-art(SOTA) performance, validating the effectiveness of our approach.
@@ I INTRODUCTION
Models based on Transformer and CNN have been widely applied in the field of transparent and reflective objects depth completion. To the best of our knowledge, however, there are currently no reported applications of the Mamba architecture in this domain. To address this long-standing gap, we conduct our research to explore the potential of the Mamba architecture in this context, aiming to open new research avenues for the community.
@@
  • We conduct extensive evaluations on multiple public datasets for transparent and reflective objects depth completion, achieving state-of-the-art(SOTA) performance.
@@ II Related Works
II-A Single-View Depth Completion
@@
Depth completion, as a fundamental task in computer vision, typically uses raw depth and RGB images to recover missing depth values in depth maps[13][14][15][16]. Due to the convenience of single-view acquisition and rapid data capture, depth completion for transparent and reflective objects under single view settings has become a research hotspot. Before end-to-end models were applied to depth completion, two approaches were adopted for transparent object depth completion. The first relies on external assumptions, such as specialized sensing devices[17] and predefined backgrounds[18]. However, these methods are unsuitable for random environments due to their dependency on external assumptions. To overcome these limitations, a second approach using global optimization algorithms emerged[19]. Studies in[20],[21]utilized sensor-captured RGB-D images as input for depth completion, eliminating the need for external assumptions. However, they rely on surface normal and edge predictions, a limitation later addressed by end-to-end models. In[22], a generative adversarial network (GAN) was employed to generate depth maps for transparent object reconstruction. To enable cross-level feature fusion and better leverage low-level depth features, [23] introduces a skip-connected fusion branch within a U-Net framework. In[24], an end-to-end U-shaped depth completion framework is proposed, combining RGB and depth images as input. [25] uses a dual-branch CNN-Transformer architecture and designs a multi-scale fusion module to integrate features at different scales. [26] designs a decoder with multiple cascaded refinement modules.
II-B Hybrid CNN-Transformer Architecture
In visual recognition tasks, Vision Transformer (ViT) models have achieved remarkable success due to their superior performance[27]. However, the multi-head self-attention (MHA) mechanism not only incurs high computational costs but also struggles to capture inductive biases for local relationships. Convolutional networks exhibit strong capabilities in capturing local features with relatively low computational cost. Consequently, integrating CNN and Transformer architectures has naturally emerged as a popular research direction, with numerous studies dedicated to leveraging the complementary strengths of both frameworks. For instance, Li et al.[28]introduced MSA at the front end of the model while employing CNNs at the backend, a design that contrasts sharply with the method in[29], where CNNs first extract local features and MSA then learns long-range dependencies. The CNN-Transformer hybrid architecture has also garnered considerable attention in the fields of medical image processing and remote sensing. For example, TransUNet[30] is the first work in medical image segmentation that integrates a U-shaped CNN with a Transformer, effectively combining local and global feature representations. In [31], novel channel-wise and spatial feature fusion modules are introduced to integrate local features extracted by the CNN branch and global context captured by the Transformer branch. Furthermore, [32] employs a dual-branch CNN to extract shallow multimodal features and leverages both self-attention and cross-attention mechanisms for deep multimodal fusion, achieving remarkable performance in the remote sensing domain.